# Modeling and computation of the effective elastic behavior of parallelogram origami metamaterials

### v2

@@ 5.3 Pinching deformations
Plots of the normalized force versus the pinching strain are shown in Fig. 9(a) for the Miura and Fig. 10(a) for the Eggbox. The force is calculated numerically using a variational method detailed in Appendix E; it is scaled by the folding modulus $d_{3}$ to give the normalized force in the plots. We also reproduce from Section 5.1 the pure mechanism results on these plots, as well as the case where the boundary condition $\mathbf{y}_{h}=\mathbf{A}(\overline{\theta})\mathbf{x}$ is applied to the full left and right parts of the boundary of $\partial\Omega$. Figs. 9 and 10(b) display the variation of the angle field versus the pinching strain in each simulation and for each boundary condition, while Figs. 9 and 10(c-e) highlight representative examples of the pinching simulations at large strains. Finally, Table 1 compares the different energy contributions for the simulations at $20\%$ strain intervals.
@@ 5.4 Bending deformations under transverse loads
One nice feature of capturing the trends in a simplified setting is that it enables design optimization without taxing a supercomputer. As a quick example, we can replace the expressions for $\lambda_{i}(\theta)$ in (50) and (51) with $\lambda_{i}(\theta+\psi)$ for an angle field $\psi$, which alters the reference folded state of the unit cells, and then optimize the stiffness as a function of $\psi$ using our simplified model. In the Miura case, the stiffness is unsurprisingly monotonic — it achieves its minimum at the fully flat state and becomes infinite at the fully folded state. The Eggbox too has a stiffness that approaches infinity at the fully folded state. However, it is most compliant at the nontrivial value $\psi\approx 0.21$. Thus, sharply corrugated patterns have high stiffness, while extremal compliance is a more delicate matter, especially in non-Euclidean cases like the Eggbox.
We end by pointing out an intriguing connection between pinching and linearized bending in these metamaterials: Both situations are well characterized by underlying second-order PDEs whose “PDE type”, and thus its basic qualitative properties, is completely dictated by the pattern’s auxeticity. Specifically, the actuation field in the pinching simulations is well approximated by the PDE for $\theta_{\text{g}}$ in (67), while the out-of-plane displacement $pw$ under small transversed load is constrained by (70). Both PDEs are elliptic for Miura Origami (or really any auxetic parallelogram origami) and hyperbolic for Eggbox (any pattern that is not auxetic). More generally, they emerge as different analytical case-studies of the geometric constraints $\mathbf{I}(\mathbf{y})=\mathbf{A}^{T}(\theta)\mathbf{A}(\theta)$ and $\big{(}\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta)\big{)}\tilde{\mathbf{v}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{v}}_{0}=-\big{(}\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta)\big{)}\tilde{\mathbf{u}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{u}}_{0}$ that underly our modeling framework. In fact, we anticipate that any such case-study of these constraints will lead to PDE compatibility conditions that illuminate an “elliptic $=$ auxetic” versus “hyperbolic $=$ non-auxetic” dichotomy in the response, although we leave this point, to be examined more carefully, for another time.
6 Conclusions
@@
Challenges and opportunities remain. It would certainly be appealing to compare our model and numerics to experiments, both qualitatively and quantitatively. We anticipate that, with proper tuning of the parameters, our approach will capture the gross-shape change observed in experiments across designs and loading condition. However, matching the loading curves may be more challenging. Potential improvements are readily available. We took a purposefully simple approach to modeling the higher-order effects in the theory. It is possible, for instance, to replace the $\int_{\Omega}(d_{1}L_{\Omega}^{2}|\nabla\nabla\mathbf{y}|^{2}+d_{2}L_{\Omega}^{2}|\nabla\theta|^{2})\mathop{}\!\mathrm{d}x$ term in the energy with the more complicated but asymptotically correct effective plate energy in (17). A mathematical challenge is to prove the existence of minimizers in this setting — the proof in Theorem 1 makes key use of the fact that the $L^{2}$-norm of $\nabla\nabla\mathbf{y}$ is controlled by the elastic energy, and will need to be adapted. Another natural modification is to replace the actuation energy $\int_{\Omega}d_{3}|\theta|^{2}\mathop{}\!\mathrm{d}x$ with a weighted average $\int_{\Omega}d_{3}\sum_{i=1,\ldots,4}\lambda_{i}|\gamma_{i}(\theta)|^{2}\mathop{}\!\mathrm{d}x$, where $\gamma_{i}\colon(\theta^{-},\theta^{+})\rightarrow\mathbb{R}$ parameterize the change in the four distinct dihedral angles of the unit cell and $\lambda_{i}$ are the weights. [^13] Presumably, each $\lambda_{i}$ should be determined geometrically by the corresponding creases perimeter or some notion of its influence area. Whether such modifications are “worth the bang for your buck” will require synergistic modeling and experimental efforts going forward.
@@ Appendix A Proof of the existence of minimizers for the constitutive model
Step 1. Extensions of the energy densities. We find it useful to express the internal energy $E_{\text{int}}(\mathbf{y},\theta)$ defined by (19-24) as
@@
for some constant $C_{2}>0$ that depends only on $\Omega$, where $\|\overline{\mathbf{y}}\|_{L^{2}(\Gamma_{\text{d}})}^{2}<\infty$ since $\overline{\mathbf{y}}\in H^{1/2}(\Gamma_{\text{d}},\mathbb{R}^{3})$. Combining the inequalities in (101-103) gives
@@ Appendix B On the derivation of the governing equations
where the term proportional to $\kappa$ in the first line vanishes since $\mathcal{H}$ is symmetric ($[\mathcal{H}]_{i\alpha\beta}=[\mathcal{H}]_{i\beta\alpha}$) and $(\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp}-\mathbf{n}\otimes\mathbf{n})$ is skew symmetric. We complete the derivation by substituting the results of (115-117) into (LABEL:eq:importForStrongForm1) and using that $\operatorname{div}\mathcal{H}=\nabla\mathcal{H}\colon\mathbf{I}$ and $\partial_{\mathbf{n}}\mathbf{w}=(\nabla\mathbf{w})\mathbf{n}$.
@@ Appendix E Computation of consistent reaction forces
[…]

### v3

@@ 5.3 Pinching deformations
Plots of the normalized force versus the pinching strain are shown in Fig. 9(a) for the Miura and Fig. 10(a) for the Eggbox. The force is calculated numerically using a variational method detailed in Appendix E; it is scaled by the folding modulus $d_{3}$ to give the normalized force in the plots. We also reproduce from Section 5.1 the pure mechanism results on these plots, as well as the case where the boundary condition $\mathbf{y}_{h}=\mathbf{A}(\overline{\theta})\mathbf{x}$ is prescribed on the full left and right parts of the boundary of $\partial\Omega$. Figs. 9 and 10(b) display the variation of the angle field versus the pinching strain in each simulation and for each boundary condition, while Figs. 9 and 10(c-e) highlight representative examples of the pinching simulations at large strains. Finally, Table 1 compares the different energy contributions for the simulations at $20\%$ strain intervals.
@@ 5.4 Bending deformations under transverse loads
One nice feature of capturing the trends in a simplified setting is that it enables design optimization without taxing a supercomputer. As a quick example, we can replace the expressions for $\lambda_{i}(\theta)$ in (53) and (54) with $\lambda_{i}(\theta+\psi)$ for an angle field $\psi$, which alters the reference folded state of the unit cells, and then optimize the stiffness as a function of $\psi$ using our simplified model. In the Miura case, the stiffness is unsurprisingly monotonic — it achieves its minimum at the fully flat state and becomes infinite at the fully folded state. The Eggbox too has a stiffness that approaches infinity at the fully folded state. However, it is most compliant at the nontrivial value $\psi\approx 0.21$.
We end by pointing out an intriguing connection between pinching and linearized bending in these metamaterials: Both situations are well characterized by underlying second-order PDEs whose “PDE type”, and thus its basic qualitative properties, is completely dictated by the pattern’s auxeticity. Specifically, the actuation field in the pinching simulations is well approximated by the PDE for $\theta_{\text{g}}$ in (69), while the out-of-plane displacement $pw$ under small transversed load is constrained by (72). Both PDEs are elliptic for Miura Origami (or really any auxetic parallelogram origami) and hyperbolic for Eggbox (any pattern that is not auxetic). More generally, they emerge as different analytical case-studies of the geometric constraints $\mathbf{I}(\mathbf{y})=\mathbf{A}^{T}(\theta)\mathbf{A}(\theta)$ and $\big{(}\mathbf{u}(\theta)\cdot\mathbf{u}^{\prime}(\theta)\big{)}\tilde{\mathbf{v}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{v}}_{0}=-\big{(}\mathbf{v}(\theta)\cdot\mathbf{v}^{\prime}(\theta)\big{)}\tilde{\mathbf{u}}_{0}\cdot\mathbf{II}(\mathbf{y})\tilde{\mathbf{u}}_{0}$ that underly our modeling framework. In fact, we anticipate that any such case-study of these constraints will lead to PDE compatibility conditions that illuminate an “elliptic $=$ auxetic” versus “hyperbolic $=$ non-auxetic” dichotomy in the response, although we leave this point to be examined more carefully for another time.
6 Discussion
@@
Figure 12: Comparing the model to an experiment. (a) A Miura origami made of perforated and folded paper is pinched at $\approx 63\%$ engineering strain. (b) Simulations of the effective model, with the parameters tuned to fit the experiment. (c) The corresponding origami deformation of the simulation.
Challenges and opportunities remain. It would certainly be appealing to compare our model and numerics to experiments, both qualitatively and quantitatively. We anticipate that, with proper tuning of the parameters, our approach will capture the gross-shape change observed in experiments across designs and loading conditions. As a demonstration, Fig. 12 highlights the model’s ability to capture qualitative features of the pinched Miura origami sample in Fig. 1(a). In brief, the experimental sample has panels of length $1$ cm and sector angle $\pi/3$ (see Fig. 5(a)), and its natural partly folded reference configuration has a cell width $\lambda_{2}(0)\approx 1.25$ cm (see Fig. 5(b)). Thus, for the simulations, we set $\lambda_{1,2}(\theta)=\lambda_{1,2}^{\text{(MO)}}(0.35\pi+\theta)$ from (53), $\Omega=(0,\lambda_{1}(0))\times(0,\lambda_{2}(0))$ for $\lambda_{1}(0)\approx 1.19$ and $\lambda_{2}(0)\approx 1.25$, and $L_{\Omega}=1.2$. The sample is pinched to $\approx 63\%$ engineering compressive strain by the pins, which are roughly $1$ unit cell in length. So we used the pinching boundary conditions in Section 5.3 for the simulations, with $\Gamma_{L,R}$ making up $10\%$ of the boundaries and with $\varepsilon_{1}(\overline{\theta})=0.63$ (see (LABEL:eq:normForce)). The simulations are performed for $c_{1}=5$ and $d_{3}=10^{-1}$ for simplicity [^14] One really needs access to experimental loading curves to choose these parameters more systematically., and for various values of $d_{1}=d_{2}$ to find one that compares well with the experiment. Fig 12(b) shows our best match, which is for $d_{1}=d_{2}=4\times 10^{-3}$; Fig. 12(c) shows the corresponding origami deformation. All told, the agreement is excellent.
This success notwithstanding, matching loading curves to experiments may be more challenging. Potential improvements are readily available. We took a purposefully simple approach to modeling the higher-order effects in the theory. It is possible, for instance, to replace the $\int_{\Omega}(d_{1}L_{\Omega}^{2}|\nabla\nabla\mathbf{y}|^{2}+d_{2}L_{\Omega}^{2}|\nabla\theta|^{2})\mathop{}\!\mathrm{d}x$ term in the energy with the more complicated but asymptotically correct effective plate energy in (19). A mathematical challenge is to prove the existence of minimizers in this setting — the proof in Theorem 1 makes key use of the fact that the $L^{2}$-norm of $\nabla\nabla\mathbf{y}$ is controlled by the elastic energy, and will need to be adapted. Another natural modification is to replace the actuation energy $\int_{\Omega}d_{3}|\theta|^{2}\mathop{}\!\mathrm{d}x$ with a weighted average $\int_{\Omega}d_{3}\sum_{i=1,\ldots,4}\lambda_{i}|\gamma_{i}(\theta)|^{2}\mathop{}\!\mathrm{d}x$, where $\gamma_{i}\colon(\theta^{-},\theta^{+})\rightarrow\mathbb{R}$ parameterize the change in the four distinct dihedral angles of the unit cell and $\lambda_{i}$ are the weights. [^15] Presumably, each $\lambda_{i}$ should be determined geometrically by the corresponding creases perimeter or some notion of its influence area. Whether such modifications are “worth the bang for your buck” will require synergistic modeling and experimental efforts going forward.
@@ Appendix A Proof of the existence of minimizers for the constitutive model
Step 1. Extensions of the energy densities. We find it useful to express the internal energy $E_{\text{int}}(\mathbf{y},\theta)$ defined by (21-27) as
@@
for some constant $C_{2}>0$ that depends only on $\Omega$, where $\|\overline{\mathbf{y}}\|_{L^{2}(\Gamma_{\text{d}})}^{2}<\infty$ since $\overline{\mathbf{y}}\in H^{1/2}(\Gamma_{\text{d}},\mathbb{R}^{3})$. Combining the inequalities in (103-105) gives
@@ Appendix B On the derivation of the governing equations
where the term proportional to $\kappa$ in the first line vanishes since $\mathcal{H}$ is symmetric ($[\mathcal{H}]_{i\alpha\beta}=[\mathcal{H}]_{i\beta\alpha}$) and $(\mathbf{n}^{\perp}\otimes\mathbf{n}^{\perp}-\mathbf{n}\otimes\mathbf{n})$ is skew symmetric. We complete the derivation by substituting the results of (117-119) into (LABEL:eq:importForStrongForm1) and using that $\operatorname{div}\mathcal{H}=\nabla\mathcal{H}\colon\mathbf{I}$ and $\partial_{\mathbf{n}}\mathbf{w}=(\nabla\mathbf{w})\mathbf{n}$.
@@ Appendix E Computation of consistent reaction forces
The numerical formulation possess a modified weak formulation, but the same strategy applies. Once we have a solution $(\mathbf{y}_{h},\theta_{h})$ in the sense that $\mathscr{A}^{h}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h}))=\mathscr{B}_{0}(\mathbf{w}_{h})$ for all $(\mathbf{w}_{h},\theta_{h})\in V_{0}^{h}$ (see Section 4.2), we then test this solution with a $(\mathbf{w}_{h,\text{f}},0)$ to reveal the reaction forces. For the pinching simulations, we choose a test function such that $\mathbf{w}_{h,\text{f}}=\mathbf{0}$ on $\Gamma_{L}$ and $\mathbf{w}_{h,\text{f}}=\mathbf{e}_{1}$ on $\Gamma_{R}$, analogous to the prior choice. The normalized compressive force plotted for these simulations is thus
$f_{1}=-\frac{1}{d_{3}}\Big{(}\mathscr{A}^{h}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h,\text{f}},0))-\mathscr{B}_{0}(\mathbf{w}_{h,\text{f}})\Big{)}.$ (132)
Appendix F Constructing origami deformations from the effective fields
In Section 5, we plot several examples of (approximate) origami deformations. Here we explain how to construct such deformations using only the effective fields $(\mathbf{y},\theta)$ as input. Along the way, we make liberal use of the notation in Section 2.1 for the design of a parallelogram origami and its perfect mechanism.
The general idea is sketched in Fig. 13. Start by recognizing that a reference parallelogram origami unit cell of characteristic length $\sim\ell$ is defined by nine vertices:
$\mathbf{x}_{0}^{\ell}=\mathbf{0},$ (133)
$\mathbf{x}_{k}^{\ell}=\ell\mathbf{t}^{r}_{k},\quad k=1,\ldots,4,$
$\mathbf{x}_{5}^{\ell}=\ell(\mathbf{t}^{r}_{1}+\mathbf{t}_{2}^{r}),\quad\mathbf{x}_{6}^{\ell}=\ell(\mathbf{t}^{r}_{2}+\mathbf{t}_{3}^{r}),\quad\mathbf{x}_{7}^{\ell}=\ell(\mathbf{t}^{r}_{3}+\mathbf{t}_{4}^{r}),\quad\mathbf{x}_{8}^{\ell}=\ell(\mathbf{t}^{r}_{4}+\mathbf{t}_{1}^{r}).$
Now fix a 2D domain $\Omega$. An overall parallelogram origami pattern can be built systematically on top of $\Omega$ by first defining the index set on $\Omega$
$I^{\ell}=\big{\{}(m,n)\in\mathbb{Z}^{2}\colon m\tilde{\mathbf{u}}_{0}+n\tilde{\mathbf{v}}_{0}\in\Omega\big{\}}$ (134)
and then he collection of vertices
$\Omega^{\ell}:=\{\mathbf{x}_{k}^{\ell}+m\ell\mathbf{u}_{0}+n\ell\mathbf{v}_{0}\colon k=0,\ldots,9\text{ and }(m,n)\in I^{\ell}\}\subset\mathbb{R}^{3},$ (135)
which is the desired parallelogram origami pattern. A stylized version of this construction is illustrated in Fig 13(a). Unlike the figure, $\Omega^{\ell}$ is a 3D collection of points overlaid on a 2D domain $\Omega$.
Figure 13: Stylized sketch and notation for constructing origami deformations. (a) A sketch of an origami pattern overlaid onto a domain $\Omega$. Each square of characteristic length $\sim\ell$ represents a panel; four squares represent a cell; $\bar{\mathbf{x}}$ labels the unit cells. (b) Sketch illustrating $\bar{\mathbf{x}}$-cell and its neighbors to the right and above. (c) Cell-based anastz is constructed for each $\bar{\mathbf{x}}$-cell. It generically produces gaps between neighboring cells. (d) The gaps are closed by averaging.
Our goal is to deform the vertices in $\Omega^{\ell}$ as approximate origami, characterized by the effective fields $(\mathbf{y},\theta)$. For this purpose, we assume that $(\mathbf{y},\theta)\colon\Omega\rightarrow\mathbb{R}^{3}\times(\theta^{-},\theta^{+})$ either solve the equilibrium equations in (43) (analytically or numerically) or satisfy the purely geometric constraints in (63). We also assume that $|\partial_{1}\mathbf{y}\times\partial_{2}\mathbf{y}|>0$ to avoid singular surfaces. Our construction then employs the following rotation field built from $\mathbf{y}$. Observe that $\mathbf{F}:=(\nabla\mathbf{y},\mathbf{n}(\mathbf{y}))$ for $\mathbf{n}(\mathbf{y})$ in (14) is a field on $\mathbb{R}^{3\times 3}$ with positive determinant. Thus, the Polar decomposition theorem furnishes the stretch tensor field $\mathbf{U}:=\sqrt{\mathbf{F}^{T}\mathbf{F}}$. The desired rotation field on $\Omega$ is then $\mathbf{R}:=\mathbf{F}\mathbf{U}^{-1}$.
We build the origami deformation from the fields $\mathbf{y},\theta,\mathbf{R}$ on $\Omega$ in two steps. The first step is a cell-based construction, where each unit cell deforms as an origami mechanism. Fix $\bar{\mathbf{x}}=\bar{m}\tilde{\mathbf{u}}_{0}+\bar{n}\tilde{\mathbf{v}}_{0}$ for $(\bar{m},\bar{n})\in I^{\ell}$ as illustrated in Fig. 13(a). The $\bar{\mathbf{x}}$-cell contains nine vertices
$\mathbf{x}_{k,\bar{\mathbf{x}}}^{\ell}=\begin{pmatrix}\bar{\mathbf{x}}\\ 0\end{pmatrix}+\mathbf{x}_{k}^{\ell},\quad i=0,\ldots,9.$ (136)
We deform the cell by taking each $\mathbf{x}_{k,\bar{\mathbf{x}}}^{\ell}$ to a $\mathbf{y}_{k,\bar{\mathbf{x}}}^{\ell}$ given by
$\mathbf{y}_{0,\bar{\mathbf{x}}}^{\ell}=\mathbf{y}(\bar{\mathbf{x}})$ (137)
$\mathbf{y}_{k,\bar{\mathbf{x}}}^{\ell}=\mathbf{y}(\bar{\mathbf{x}})+\ell\mathbf{R}(\bar{\mathbf{x}})\mathbf{t}_{k}(\theta(\bar{\mathbf{x}})),\quad k=1,\ldots,4,$
$\mathbf{y}_{5,\bar{\mathbf{x}}}^{\ell}=\mathbf{y}(\bar{\mathbf{x}})+\ell\mathbf{R}(\bar{\mathbf{x}})\big{[}\mathbf{t}_{1}(\theta(\bar{\mathbf{x}}))+\mathbf{t}_{2}(\theta(\bar{\mathbf{x}}))\big{]},$
$\mathbf{y}_{6,\bar{\mathbf{x}}}^{\ell}=\mathbf{y}(\bar{\mathbf{x}})+\ell\mathbf{R}(\bar{\mathbf{x}})\big{[}\mathbf{t}_{2}(\theta(\bar{\mathbf{x}}))+\mathbf{t}_{3}(\theta(\bar{\mathbf{x}}))\big{]},$
$\mathbf{y}_{7,\bar{\mathbf{x}}}^{\ell}=\mathbf{y}(\bar{\mathbf{x}})+\ell\mathbf{R}(\bar{\mathbf{x}})\big{[}\mathbf{t}_{3}(\theta(\bar{\mathbf{x}}))+\mathbf{t}_{4}(\theta(\bar{\mathbf{x}}))\big{]},$
$\mathbf{y}_{8,\bar{\mathbf{x}}}^{\ell}=\mathbf{y}(\bar{\mathbf{x}})+\ell\mathbf{R}(\bar{\mathbf{x}})\big{[}\mathbf{t}_{4}(\theta(\bar{\mathbf{x}}))+\mathbf{t}_{1}(\theta(\bar{\mathbf{x}}))\big{]}.$
Applying these formulas to all the $\bar{\mathbf{x}}$-cells defined by $\Omega^{\ell}$ produces a cell-based ansatz $\{\mathbf{y}^{\ell}_{k,\bar{\mathbf{x}}}\}$. Each cell deforms as a perfect mechanism informed by the fields $\mathbf{y},\theta$ and $\mathbf{R}$. However, intercell gaps generically arise since we do not enforce any compatibility conditions at the cell boundaries. Fig. 13(b) and (c) qualitatively sketches this ansatz and its gaps.
The final step of the construction is to average the gaps, and thus replace the cell-based ansatz $\{\mathbf{y}_{k,\bar{\mathbf{x}}}^{\ell}\}$ by a vertex based one $\{\mathbf{y}_{i,j}^{\ell}\}$. This is basically an accounting problem, involving a number of distinct cases. We detail one of them to explain the main idea. Let $(i^{\prime},j^{\prime})\in\mathbb{Z}^{2}$ index the one vertex that belongs to all four cells in Fig. 13(b), as shown in Fig. 13(d). The cell based construction in Fig. 13(c) maps this point to (potentially) four distinct points by via the cell formulas in (LABEL:eq:cellBaased). We now simply average those points to close the gaps
$\mathbf{y}_{i^{\prime},j^{\prime}}^{\ell}=\frac{1}{4}\big{(}\mathbf{y}_{5,\bar{\mathbf{x}}}^{\ell}+\mathbf{y}_{6,\bar{\mathbf{x}}+\ell\tilde{\mathbf{u}}_{0}}^{\ell}+\mathbf{y}_{8,\bar{\mathbf{x}}+\ell\tilde{\mathbf{v}}_{0}}^{\ell}+\mathbf{y}_{7,\bar{\mathbf{x}}+\ell(\tilde{\mathbf{u}}_{0}+\tilde{\mathbf{v}}_{0})}^{\ell}\big{)}$ (138)
A similar strategy is employed to close all the gaps in the pattern, completing the construction.
Note that we have added panel strain by this averaging step. However, the strain should be quite small. Basically, it should be imperceptible to the naked eye, so long as the cell size $\ell$ is $\ll$ than the characteristic length of the domain $L_{\Omega}$ and the fields $(\mathbf{y},\theta)$ satisfy or approximately satisfy (63). For those interested, a rigorous justification of this statement is one of the main topics of our prior work [64].